# A Lie algebraic pattern behind logarithmic CFTs

### v2

@@ 4.2. The non-super case (1)
Let us consider our case (1). By the injectivity of the Miura map [Gen17, Lemma 5.4] and the consideration in [FKRW95, Remark 4.1], we have
@@
where $b_{i}$ satisfies $(b_{i},b_{j})=(k+h^{\vee})(\alpha_{i},\alpha_{j})$. In particular, Fock module $M(\mu)$ is also a $\mathbf{W}^{k}(\mathfrak{g})$-module. By comparing the conformal vector (2.10) and [ArF19, (3.22)], for $k=\frac{1}{p}-h^{\vee}$ and $b_{i}=-\frac{1}{\sqrt{p}}\alpha_{i}$, we have
@@
On the other hands, by [ArF19, Section 4.3], we have
$\operatorname{ch}_{q}T^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}=\operatorname{ch}_{q}T^{\frac{1}{p}}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}=\operatorname{ch}_{q}\mathcal{W}_{-\alpha+\lambda}$ (4.4)
@@
$\mathcal{W}_{0}\simeq\mathbf{W}^{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})\simeq\mathbf{W}_{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})\simeq\mathbf{W}^{\frac{1}{p}-h^{\vee}}(\mathfrak{g})\simeq\mathbf{W}_{\frac{1}{p}-h^{\vee}}(\mathfrak{g}).$ (4.5)
@@
$k=\tfrac{1}{p}-h^{\vee},\ \ \check{k}=m-{}^{L}h^{\vee}.$ (4.6)
On the other hand, in the same manner as [Sug22, Lemma 3.14, 3.19], we have the following.
@@
Our aim is to apply the discussion in Section 2.2 (see the last two paragraphs). First let us give a Kazhdan-Lusztig type decomposition in Lemma 2.6(3). Unless otherwise noted, all symbols (e.g., $Q$, $\alpha_{i}$,…) are the ones of $\mathfrak{g}$. Denote $h_{\mu}$ and $\check{h}_{\mu}$ the top conformal weight of $\mathbf{M}_{k}(\chi_{\mu})$ and $\check{\mathbf{M}}_{\check{k}}(\chi_{\mu})$, respectively (see [ACL19, (30)]). By (2.6), for any $\beta\in(\sqrt{p}Q)^{\ast}$ (in other words, $\sqrt{p}\beta\in P_{{}^{L}\mathfrak{g}}$), we have
$\Delta_{\beta}=h_{-\frac{1}{\sqrt{p}}\beta}=\check{h}_{\sqrt{p}\beta}.$ (4.7)
@@
Let us recall the notation in [KT00]. In our case (i.e., $\widehat{{}^{L}\mathfrak{g}}$), we have
@@
${\mathcal{C}}^{+}_{\widehat{{}^{L}\mathfrak{g}}}=\{\nu\ |\ (\gamma^{\vee},\nu+\hat{\rho}_{{}^{L}\mathfrak{g}})_{{}^{L}Q}\geq 0\ \text{for any $\gamma\in\Delta^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}}$}\}=\{\nu\ |\ (\gamma^{\vee},\nu+\rho^{\vee}+{}^{L}h^{\vee}\Lambda_{0,{}^{L}\mathfrak{g}})_{{}^{L}Q}\geq 0\ \text{for any $\gamma^{\vee}\in(\Delta^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}})^{\vee}$}\}.$
By [KT00, Lemma 2.10], there exists unique $\sigma t_{r^{\vee}\beta}\in\hat{W}_{{}^{L}\mathfrak{g}}=W\ltimes r^{\vee}Q$ such that
@@
[…]
We show that if $\lambda$ satisfies the condition Lemma 2.9(1), i.e. $(p\lambda_{\bullet}+\rho^{\vee},\theta_{s})\leq m$, such $\sigma t_{r^{\vee}\beta}$ is independent of the choice of $\lambda_{\bullet}$. Let us consider the case $\mathfrak{g}=B_{r}$ (other cases are similar). Set $\beta-(\alpha+\lambda^{\bullet}+\rho)=\sum_{i\in\Pi}n_{i}\varpi_{i}$ ($n_{i}\in\mathbb{Z}$). For $\alpha_{+}\in\Delta_{+}$, if there exists, denote $a_{\pm\alpha_{+}}\in\Delta_{+}$ such that $\sigma^{-1}(a_{\pm\alpha_{+}})=\pm\alpha_{+}$. For any $\alpha_{+}\in\Delta_{+}$, one of $a_{\alpha_{+}}$ or $a_{-\alpha_{+}}$ always exists. Set $I_{\pm}=\{i\ |\ \exists a_{\pm\alpha_{i}}\}$. Clearly, we have $I=I_{+}\sqcup I_{-}$. For $i\in I_{+}$, we have
@@
and thus $n_{i}=0$ for $i\in I_{+}$. Similarly, we have $n_{i}=-1$ for $i\in I_{-}$, and thus $\sum_{i\in I}n_{i}\varpi_{i}=-\sum_{i\in I_{-}}\varpi_{i}$. Next, let us consider $a_{\pm\theta_{s}}$. If $a_{\theta_{s}}$ exists, then we have $I_{-}=\phi$. On the other hand, if $a_{-\theta_{s}}$ exists, then we have $I_{-}=\{r\}$. Each of these cases corresponds to the case $\alpha+\lambda^{\bullet}+\rho\in Q$ or not, which is independent of the choice of $\lambda_{\bullet}$. Therefore, we can use the notation $y_{\alpha,\lambda^{\bullet}}\in\hat{W}$ for the inverse of the unique element $\sigma t_{\beta}$ in above discussion. By [KT00, Theorem 1.1], we obtain the following:
@@
$\mu_{\lambda}=y_{\alpha,\lambda^{\bullet}}^{-1}\circ_{\widehat{{}^{L}\mathfrak{g}}}(-p(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\check{k}\Lambda_{0,{}^{L}\mathfrak{g}}).$ (4.8)
Furthermore, by applying the (exact) $+$-reduction functor (see [Ar07, Theorem 7.7.1]), we have
@@
respectively. Then the conditions Lemma 2.6(3),(4) are already checked. On the other hand, the condition Lemma 2.6(1),(2) follows from (4.7) and [ACL19, (29)]. Then by Lemma 2.6 and Lemma 4.1, $\mathcal{W}_{-\alpha+\lambda}$ is simple as $\mathcal{W}_{0}$-module. In particular, each $\mathcal{W}_{-\alpha+\lambda}$ is a CFT-type $\mathcal{W}_{0}$-module. Finally, by applying Lemma 2.8, the simplicity of $H^{0}(G\times_{B}V_{\lambda})$ also holds. In conclusion, we have the following:
@@
Let us consider the non-super case (1), and set $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$. Then, for (4.6) and $m\geq{}^{L}h^{\vee}-1$, we have the VOA-isomorphism
@@
and each $\mathcal{W}_{-\alpha+\lambda^{\bullet}}$ is simple as $\mathcal{W}_{0}$-module ($\alpha\in P_{+}\cap Q$). More generally, for $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
[…]
@@ 4.3. The super case (2)
${k}=\tfrac{1}{2(2m-1)}-(r+\tfrac{1}{2}),\ \ \check{k}=m-r-1,\ \ h^{\vee}=h^{\vee}_{\mathfrak{osp}(1|2r)}=r+\tfrac{1}{2}.$ (4.9)
Then $14(k+h^{\vee})(\check{k}+h^{\vee})=$ 1 and by [Gen17, Theorem 6.3] and Corollary 3.6, we have
$\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))$ (4.10)
@@
We aim to apply the discussion in Section 2.2 (see the last two paragraphs). First let us consider Lemma 2.6 in our case ($\lambda_{0}=0$, $\lambda_{1}=\lambda$ such that $\lambda^{\bullet}=0$ and $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$). The condition Lemma 2.6(4) is already checked in the above discussion. By [KRW78] and the results in Section 2.3, we have
@@
Lastly, we want to check the Lemma 2.6(3). However, to the authors’ knowledge, the Kazhdan-Lusztig type decomposition as [KT00, Ar07] has not been proven in our case $\mathfrak{osp}(1|2r)$. In this paper, we will prove the Kazhdan-Lusztig type decomposition (and Lemma 2.6(3)) in the case of twisted affine Lie algebra of type $A^{(2)}_{2r}$, and prove the simplicity theorem for super-case (2) under the assumption that the characters in these two types are the same under the coordinate change. Let us recall the data in $\hat{\mathfrak{g}}^{\prime}=A_{2r}^{(2)}$:  [^10] Note that here we use the second description in [Car05]. On the other hand, [Kac90, Wak01] uses the first description of [Car05]. The classical part is different for these two descriptions: For the first one (resp. second one), it is type $C_{n}$ (resp. type $B_{n}$).
@@
where all classical terms are of type $B_{n}$ (see [Kac90, Section 6]), and $\Lambda_{0}^{c}=\frac{1}{2}\Lambda_{0}$. By [KT00, Lemma 2.10], there exists unique $\sigma t_{\beta}\in\hat{W}_{\hat{\mathfrak{g}}^{\prime}}=W\ltimes Q$ such that
@@
By combining it with [KT00, Theorem 1.1], we obtain the Kazhdan-Lusztig type decomposition in $A_{2r}^{(2)}$ case.
@@
Let us relate type $A_{2r}^{(2)}$ and $\mathfrak{osp}(1|2r)$. Because the second description of $A_{2n}^{(2)}$ in [Car05] has the same Dynkin diagram as the one of type $B(0,r)=\mathfrak{osp}(1|2r)$, replacing a black node with a white node, these two types are closely related. In particular, according to [KW01, Section 9.5], the character formulas for the modules of affine Lie (super)algebras of type $A_{2r}^{(2)}$ and of type $B(0,r)$ are the same up to coordinate change. In fact, choosing the coordinate of $h\in\hat{\mathfrak{h}}$ of twisted affine Kac-Moody Lie algebra of type $A_{2r}^{(2)}$:
@@
$\operatorname{ch}[\hat{M}(A_{2r}^{(2)})](h)=\operatorname{ch}[\hat{M}(B(0,r)](t(h)).$ (4.11)
Meanwhile, for type $A_{2n}^{(2)}$, the Kazhdan-Lusztig type decomposition was already proved in Lemma 4.5 above. Therefore, we also have the same Kazhdan-Lusztig type decomposition in $\mathfrak{osp}(1|2r)$-side.
To apply Lemma 2.6(3) to our super case (2), one must move from affine to the principal $W$-algebra. We denote the category $\mathcal{O}$ for $L_{\kappa}(\mathfrak{osp}(1|2r))$ by $\mathcal{O}_{\kappa}$ ($\kappa\in\mathbb{C}$). Let $\mathcal{O}_{\kappa}^{[\lambda]}$ be the full subcategory of $\mathcal{O}_{\kappa}$ whose objects have their local composition factors isomorphic to $\hat{L}_{\kappa}(w\circ\lambda)$. The exactness of the $+$-reduction $H^{0}_{\mathrm{DS},+}(\cdot)$ is already proved in [Ar07] (and thus the Kazhdan-Lusztig type character formula of $\mathbf{W}^{\kappa}(\mathfrak{g})$-module is derived from [KT00, Theorem 1.1]), but in our case it has not known yet to the authors’ knowledge.
@@
For $\check{k}=m-r-1$ and $\lambda\in\Lambda$ $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$, the $+$-reduction functor $H^{0}_{\mathrm{DS},+}(\cdot)$ defines an exact functor from $\mathcal{O}_{\check{k}}^{[\lambda]}$ to $\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))\text{-$\mathrm{mod}$}$.
If we assume the above conjecture, then we have the Kazhdan-Lusztig type decomposition in our case $\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$ and the last condition Lemma 2.6(3) is also satisfied. The remaining discussion is the same as Section 4.2, and the conclusion is as follows:
@@
[…]

### v3

@@ 4.2. The non-super case (1)
Let us consider our case (1). By the injectivity of the Miura map [28, Lemma 5.4] and the consideration in [26, Remark 4.1], we have
@@
where $b_{i}$ satisfies $(b_{i},b_{j})=(k+h^{\vee})(\alpha_{i},\alpha_{j})$. In particular, Fock module $M(\mu)$ is also a $\mathbf{W}^{k}(\mathfrak{g})$-module. By comparing the conformal vector (2.10) and [5, (3.22)], for $k=\frac{1}{p}-h^{\vee}$ and $b_{i}=-\frac{1}{\sqrt{p}}\alpha_{i}$, we have
@@
On the other hands, by [5, Section 4.3], we have
(4.4) $\operatorname{ch}_{q}T^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}=\operatorname{ch}_{q}T^{\frac{1}{p}}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}=\operatorname{ch}_{q}\mathcal{W}_{-\alpha+\lambda}$
@@
(4.5) $\mathcal{W}_{0}\simeq\mathbf{W}^{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})\simeq\mathbf{W}_{m-{}^{L}h^{\vee}}({}^{L}\mathfrak{g})\simeq\mathbf{W}^{\frac{1}{p}-h^{\vee}}(\mathfrak{g})\simeq\mathbf{W}_{\frac{1}{p}-h^{\vee}}(\mathfrak{g}).$
@@
(4.6) $k=\tfrac{1}{p}-h^{\vee},\ \ \check{k}=m-{}^{L}h^{\vee}.$
On the other hand, in the same manner as [46, Lemma 3.14, 3.19], we have the following.
@@
Our aim is to apply the discussion in Section 2.2 (see the last two paragraphs). First let us give a KL type decomposition in Lemma 2.6(3). Unless otherwise noted, all symbols (e.g., $Q$, $\alpha_{i}$,…) are the ones of $\mathfrak{g}$. Denote $h_{\mu}$ and $\check{h}_{\mu}$ the top conformal weight of $\mathbf{M}_{k}(\chi_{\mu})$ and $\check{\mathbf{M}}_{\check{k}}(\chi_{\mu})$, respectively (see [6, (30)]). By (2.6), for any $\beta\in(\sqrt{p}Q)^{\ast}$ (in other words, $\sqrt{p}\beta\in P_{{}^{L}\mathfrak{g}}$), we have
(4.7) $\Delta_{\beta}=h_{-\frac{1}{\sqrt{p}}\beta}=\check{h}_{\sqrt{p}\beta}.$
@@
Let us recall the notation in [37]. In our case (i.e., $\widehat{{}^{L}\mathfrak{g}}$), we have
@@
$\begin{aligned} {\mathcal{C}}^{+}_{\widehat{{}^{L}\mathfrak{g}}}&=\{\nu\ |\ (\gamma^{\vee},\nu+\hat{\rho}_{{}^{L}\mathfrak{g}})_{{}^{L}Q}\geq 0\ \text{for any $\gamma\in\Delta^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}}$}\}\\ &=\{\nu\ |\ (\gamma^{\vee},\nu+\rho^{\vee}+{}^{L}h^{\vee}\Lambda_{0,{}^{L}\mathfrak{g}})_{{}^{L}Q}\geq 0\ \text{for any $\gamma^{\vee}\in(\Delta^{\mathrm{re}}_{+,\widehat{{}^{L}\mathfrak{g}}})^{\vee}$}\}.\end{aligned}$
By [37, Lemma 2.10], there exists unique $\sigma t_{r^{\vee}\beta}\in\hat{W}_{{}^{L}\mathfrak{g}}=W\ltimes r^{\vee}Q$ such that
@@
$\land\$$0\leq(\sigma^{-1}(\Delta_{+}^{s}),p(\beta-(\alpha+\lambda^{\bullet}+\rho))+p\lambda_{\bullet}+\rho^{\vee})_{Q}\leq m.$
We show that if $\lambda$ satisfies the condition Lemma 2.9(1), i.e. $(p\lambda_{\bullet}+\rho^{\vee},\theta_{s})\leq m$, such $\sigma t_{r^{\vee}\beta}$ is independent of the choice of $\lambda_{\bullet}$. Let us consider the case $\mathfrak{g}=B_{r}$ (other cases are similar). Set $\beta-(\alpha+\lambda^{\bullet}+\rho)=\sum_{i\in I}n_{i}\varpi_{i}$ ($n_{i}\in\mathbb{Z}$). For $\alpha_{+}\in\Delta_{+}$, if there exists, denote $a_{\pm\alpha_{+}}\in\Delta_{+}$ such that $\sigma^{-1}(a_{\pm\alpha_{+}})=\pm\alpha_{+}$. For any $\alpha_{+}\in\Delta_{+}$, one of $a_{\alpha_{+}}$ or $a_{-\alpha_{+}}$ always exists. Set $I_{\pm}=\{i\ |\ \exists a_{\pm\alpha_{i}}\}$. Clearly, we have $I=I_{+}\sqcup I_{-}$. For $i\in I_{+}$, we have
@@
and thus $n_{i}=0$ for $i\in I_{+}$. Similarly, we have $n_{i}=-1$ for $i\in I_{-}$, and thus $\sum_{i\in I}n_{i}\varpi_{i}=-\sum_{i\in I_{-}}\varpi_{i}$. Next, let us consider $a_{\pm\theta_{s}}$. If $a_{\theta_{s}}$ exists, then we have $I_{-}=\phi$. On the other hand, if $a_{-\theta_{s}}$ exists, then we have $I_{-}=\{r\}$. Each of these cases corresponds to the case $\alpha+\lambda^{\bullet}+\rho\in Q$ or not, which is independent of the choice of $\lambda_{\bullet}$. Therefore, we can use the notation $y_{\alpha,\lambda^{\bullet}}\in\hat{W}$ for the inverse of the unique element $\sigma t_{\beta}$ in above discussion. By [37, Theorem 1.1], we obtain the following:
@@
(4.8) $\mu_{\lambda}=y_{\alpha,\lambda^{\bullet}}^{-1}\circ_{\widehat{{}^{L}\mathfrak{g}}}(-p(\alpha+\lambda^{\bullet}+\rho)+p\lambda_{\bullet}+\check{k}\Lambda_{0,{}^{L}\mathfrak{g}}).$
Furthermore, by applying the (exact) ++-reduction functor (see [10, Theorem 7.7.1]), we have
@@
respectively. Then the conditions Lemma 2.6(3),(4) are already checked. On the other hand, the condition Lemma 2.6(1),(2) follows from (4.7) and [6, (29)]. Then by Lemma 2.6 and Lemma 4.1, $\mathcal{W}_{-\alpha+\lambda}$ is simple as $\mathcal{W}_{0}$-module. In particular, each $\mathcal{W}_{-\alpha+\lambda}$ is a CFT-type $\mathcal{W}_{0}$-module. Finally, by applying Lemma 2.8, the simplicity of $H^{0}(G\times_{B}V_{\lambda})$ also holds. In conclusion, we have the following:
@@
Let us consider the non-super case (1) and $H^{0}(G\times_{B}V_{\lambda})\simeq\bigoplus_{\alpha\in P_{+}\cap Q}L_{\alpha+\lambda^{\bullet}}\otimes\mathcal{W}_{-\alpha+\lambda}$. Then, for (4.6) and $m\geq{}^{L}h^{\vee}-1$, we have the VOA-isomorphism
@@
and each $\mathcal{W}_{-\alpha+\lambda^{\bullet}}$ is simple as $\mathcal{W}_{0}$-module. More generally, for $\lambda\in\Lambda$ such that $(p\lambda_{\bullet}+\rho^{\vee},{}^{L}\theta)\leq p$,
$\mathcal{W}_{-\alpha+\lambda}$ $\simeq\check{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})\simeq\check{\mathbf{T}}^{m}_{p\lambda_{\bullet},\alpha+\lambda^{\bullet}}\ \ \text{(as $\mathbf{W}^{\check{k}}({}^{L}\mathfrak{g})$-modules)}$
$\simeq\mathbf{L}_{k}(\chi_{\alpha-\lambda})\simeq\mathbf{T}^{\frac{1}{p}}_{\alpha+\lambda^{\bullet},p\lambda_{\bullet}}\ \ \text{(as $\mathbf{W}^{k}(\mathfrak{g})$-modules)},$
@@
Remark 4.4.
It is expected that the simple modules of $H^{0}(G\times_{B}V_{0})$ are classified by $\Lambda$. Constructing simple modules for general $\lambda\in\Lambda$ seems to require the development of not only our geometric approach, but also the W-algebraic analysis of the Verma module $)\mathbf{M}_{k}(\chi_{\mu})=H^{0}_{\mathrm{DS},\check{\mu}}(\hat{M}(\hat{\lambda}$)). The latter is currently being pursued by T. Arakawa and the first author to investigate the tensor category structure of W-algebras at generic levels.
@@ 4.3. The super case (2)
(4.9) ${k}=\tfrac{1}{2(2m-1)}-(r+\tfrac{1}{2}),\ \ \check{k}=m-r-1,\ \ h^{\vee}=h^{\vee}_{\mathfrak{osp}(1|2r)}=r+\tfrac{1}{2}.$
Then $4(k+h^{\vee})(\check{k}+h^{\vee})=1$ and by [28, Theorem 6.3] and Corollary 3.6, we have
(4.10) $\mathcal{W}_{0}\simeq\mathbf{W}^{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{k}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))\simeq\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))$
@@
We aim to apply the discussion in Section 2.2 (see the last two paragraphs). First let us consider Lemma 2.6 in our case ($\lambda_{0}=0$, $\lambda_{1}=\lambda$ such that $\lambda^{\bullet}=0$ and $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$). The condition Lemma 2.6(4) is already checked in the above discussion. By [36] and the results in Section 2.3, we have
@@
Lastly, we want to check the Lemma 2.6(3). However, to the authors’ knowledge, the KL type decomposition as [37, 10] has not been proven in our case $\mathfrak{osp}(1|2r)$. In this paper, we will prove the KL type decomposition (and Lemma 2.6(3)) in the case of twisted affine Lie algebra of type $A^{(2)}_{2r}$, and prove the simplicity theorem for super-case (2) under the assumption that the characters in these two types are the same under the coordinate change. Let us recall the data in $\hat{\mathfrak{g}}^{\prime}=A_{2r}^{(2)}$: [^11] Note that here we use the second description in [12]. On the other hand, [34, 50] uses the first description of [12]. The classical part is different for these two descriptions: For the first one (resp. second one), it is type $C_{n}$ (resp. type $B_{n}$).
@@
where all classical terms are of type $B_{n}$ (see [34, Section 6]), and $\Lambda_{0}^{c}=\frac{1}{2}\Lambda_{0}$. By [37, Lemma 2.10], there exists unique $\sigma t_{\beta}\in\hat{W}_{\hat{\mathfrak{g}}^{\prime}}=W\ltimes Q$ such that
@@
By combining it with [37, Theorem 1.1], we obtain the KL type decomposition in $A_{2r}^{(2)}$ case.
@@
Let us relate type $A_{2r}^{(2)}$ and $\mathfrak{osp}(1|2r)$. Because the second description of $A_{2n}^{(2)}$ in [12] has the same Dynkin diagram as the one of type $B(0,r)=\mathfrak{osp}(1|2r)$, replacing a black node with a white node, these two types are closely related. In particular, according to [35, Section 9.5], the character formulas for the modules of affine Lie (super)algebras of type $A_{2r}^{(2)}$ and of type $B(0,r)$ are the same up to coordinate change. In fact, choosing the coordinate of $h\in\hat{\mathfrak{h}}$ of twisted affine Kac-Moody Lie algebra of type $A_{2r}^{(2)}$:
@@
(4.11) $\operatorname{ch}[\hat{M}(A_{2r}^{(2)})](h)=\operatorname{ch}[\hat{M}(B(0,r)](t(h)).$
Meanwhile, for type $A_{2n}^{(2)}$, the KL type decomposition was already proved in Lemma 4.6 above. Therefore, we also have the same KL type decomposition in $\mathfrak{osp}(1|2r)$-side.
To apply Lemma 2.6(3) to our super case (2), one must move from affine to the principal $W$-algebra. We denote the category $\mathcal{O}$ for $L_{\kappa}(\mathfrak{osp}(1|2r))$ by $\mathcal{O}_{\kappa}$ ($\kappa\in\mathbb{C}$). Let $\mathcal{O}_{\kappa}^{[\lambda]}$ be the full subcategory of $\mathcal{O}_{\kappa}$ whose objects have their local composition factors isomorphic to $\hat{L}_{\kappa}(w\circ\lambda)$. The exactness of the ++-reduction $H^{0}_{\mathrm{DS},+}(\cdot)$ is already proved in [10] (and thus the KL type character formula of $\mathbf{W}^{\kappa}(\mathfrak{g})$-module is derived from [37, Theorem 1.1]), but in our case it has not known yet to the authors’ knowledge.
@@
For $\check{k}=m-r-1$ and $\lambda\in\Lambda$ $(p\lambda_{\bullet}+\rho,{}^{L}\theta)\leq p$, the ++-reduction functor $H^{0}_{\mathrm{DS},+}(\cdot)$ defines an exact functor from $\mathcal{O}_{\check{k}}^{[\lambda]}$ to $\mathbf{W}_{\check{k}}(\mathfrak{osp}(1|2r))\text{-$\mathrm{mod}$}$.
If we assume the above conjecture, then we have the KL type decomposition in our case $\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$ and the last condition Lemma 2.6(3) is also satisfied. The remaining discussion is the same as Section 4.2, and the conclusion is as follows:
@@
$\mathcal{W}_{-\alpha+\lambda}$ $\simeq{\mathbf{L}}_{\check{k}}(\chi_{p(-\alpha+\lambda)})\ \ \text{(as $\mathbf{W}^{\check{k}}(\mathfrak{osp}(1|2r))$-modules)}$
$\simeq\mathbf{L}_{k}(\chi_{\alpha-\lambda})\ \ \text{(as $\mathbf{W}^{k}(\mathfrak{osp}(1|2r))$-modules)},$